### 14 Jan 2024 v1

# Unveiling Attacks Ahead of Exploit: adversarial Contract Detection Based on DeFi Attack Features

Abstract
DeFi incidents stemming from various smart contract vulnerabilities have culminated in financial damages exceeding 3 billion USD. The attacks causing such incidents commonly commence with the deployment of adversarial contracts, subsequently leveraging these contracts to execute adversarial transactions that exploit vulnerabilities in victim contracts. Existing defense mechanisms leverage heuristic or machine learning algorithms to detect adversarial transactions, but they face significant challenges in detecting private adversarial transactions. Namely, attackers can send adversarial transactions directly to miners, evading visibility within the blockchain network and effectively bypassing the detection.
In this paper, we propose a new direction for detecting DeFi attacks, i.e., detecting adversarial contracts instead of adversarial transactions, allowing us to proactively identify potential attack intentions, even if they employ private adversarial transactions. Specifically, we observe that most adversarial contracts follow a similar pattern, e.g., anonymous fund source, closed-source, frequent token-related function calls. Based on this observation, we build a machine learning classifier that can effectively distinguish adversarial contracts from benign ones.
We build a dataset consists of features extracted from 304 adversarial contracts and 13,000 benign contracts. Based on this dataset, we evaluate different classifiers, the results of which show that our method for identifying DeFi adversarial contracts performs exceptionally well. For example, the F1-Score for LightGBM-based classifier is 0.9434, with a remarkably low false positive rate of only 0.12%.
1 Introduction
Decentralized Finance (DeFi) has gained significant prominence and is evolving rapidly within the blockchain ecosystem in recent years, giving rise to a diverse array of DeFi applications such as decentralized exchanges (DEXs) [25, 43] and lending platforms [7, 49]. The total value locked (TVL) within these DeFi protocols exceeded \$50 billion in early 2024 [36]. Among various blockchain networks, Ethereum [56] and Ethereum Virtual Machine (EVM)-compatible chains [28, 26, 44, 45] have a thriving DeFi ecosystem, playing a pivotal role in the landscape.
At the heart of DeFi are smart contracts, which are self-executing programs deployed on blockchains, enabling trustless and transparent financial interactions without the need for intermediaries. They could be invoked by users through transactions and autonomously execute predefined actions. However, like traditional software, they are fundamentally code-based programs, making them susceptible to potential security risks. Indeed, DeFi incidents resulting from various smart contract vulnerabilities have caused over 3 billion USD in losses [62, 18, 23]. Figure 1 depicts two common attack patterns:
Refer to captionAttackerRefer to caption Victim contract(s) 1attack2profit
(a) Attacking directly.
Refer to captionAttackerRefer to caption Victim contract(s) Refer to captionAdversarial contract1deploy2callattack3profit4
(b) Attacking through adversarial contracts.
Figure 1: Two common patterns for DeFi attacks: (a) directly interacts with the victim contracts (b) interacts through adversarial contracts.
  1. 1)
    Attacking directly. The attacker directly invokes the functions of the victim contracts [11, 13]. However, in the presence of complex attacking logic involving multiple transactions, there is a potential risk of half-way failure. This situation arises when part of the attack is executed but the rest cannot proceed (e.g., due to transaction conflictions).
  2. 2)
    Attacking through adversarial contracts. The attacker first deploys an adversarial contract that contains specific attack scripts for the victim contracts, and then initiates the attack by calling the entry function of the adversarial contract[10, 12, 8, 24]. This pattern allows the attacker to execute complex attacking logic in a single transaction, hence avoid half-way failure due to the atomic nature of a transaction. Most DeFi attacks follow this pattern.
Existing defense mechanisms leverage heuristic [60, 63, 46, 58] or machine learning [37, 54] algorithms to detect adversarial transactions, but they face significant challenges in detecting private adversarial transactions. Namely, attackers can send adversarial transactions directly to miners, evading visibility within the blockchain network and effectively bypassing the detection. Based on our analysis of historical DeFi incidents, we found that the proportion of adversarial transactions sent using privacy services has exhibited a significant upward trend, rising from 33% in the first half of 2022 to a notable 71% in the latter half of 2023.
In this paper, we propose a new direction for detecting DeFi attacks, i.e., detecting adversarial contracts instead of adversarial transactions, allowing us to proactively identify potential attack intentions, even if they employ private adversarial transactions.
We first collected 304 adversarial contracts used in DeFi incidents between April 2020 and December 2023 from both the Ethereum and Binance Smart Chain (BSC). We also collected the top 13,000 contracts on Ethereum from June 2022 to June 2023, to serve as our benign contract samples. These data are made publicly available to facilitate future research in this domain.
We observe that most adversarial contracts follow a similar pattern, e.g., anonymous fund source, closed-source, frequent token-related function calls. Based on this observation, we extract six categories of features from each contract, such as attributes of contract creator, implemented functions in contract, and contract types. We then build classifiers using multiple supervised machine learning algorithms and evaluate them comprehensively. The results show that our method for identifying DeFi adversarial contracts performs exceptionally well. For example, the F1-Score for LightGBM-based classifier is 0.9434, with a remarkably low false positive rate of only 0.12%.
We summarize our contributions as follows:
  • A new direction for detecting DeFi attacks, i.e., detecting adversarial contracts instead of adversarial transactions.
  • The first dataset for adversarial contracts.
  • A comprehensive evaluation of different classifiers.
2 Background
In this section, we outline the required background to better understand our work.
2.1 Ethereum
Ethereum[56] is an open-source, decentralized blockchain platform that introduces EVM to support smart contracts and has spawned a series of EVM-compatible chains.
Account. Accounts are entities that can hold Ether and send transactions. Ethereum has two main types of accounts: External Owned Account (EOA) and Contract Account (CA). EOAs are controlled by users’ private keys. Both types of accounts have their address, balance, and nonce.
Transcation. Ethereum is often regarded as a transaction-driven state machine, where transactions are initiated by EOA, signed, and ultimately update the Ethereum network’s state. Sending transactions require Ether as transaction fees. Transactions are typically used for transferring Ether, invoking contracts, or creating contracts. In the case of contract deployment, the to field is 0. During the execution of a normal transaction, interacted contracts may trigger additional message calls, commonly known as internal transactions.
Smart Contract. From a high-level perspective, smart contracts can be viewed as instantiated objects stored on the blockchain and executed within the EVM. Due to the nature of blockchain, once deployed, the code within a smart contract cannot be changed. Smart contracts are typically written in high-level programming languages such as Solidity [35]. The source code is compiled to bytecode, which can be executed within the EVM. The EVM is a stack-based virtual machine that supports a set of Turing-complete instructions. Bytecode mainly consists of two parts, creation code and runtime code. Creation code is used for initialization (not deployed on the blockchain) and is executed only once. Runtime code is the primary code segment executed when the contract receives a message call.
2.2 DeFi
DeFi is a financial ecosystem that operates on a blockchain network. It is built upon smart contracts and aims to provide traditional financial services in a more open and transparent manner. DeFi has also introduced some unique financial concepts, such as tokens, decentralized exchange, and flashloan.
The tokens. In addition to Ether, there are numerous cryptocurrencies created by smart contracts on Ethereum, known as tokens. Tokens are the foundation of the DeFi ecosystem, supporting transactions and activities within DeFi. Tokens come in two main types: fungible tokens and non-fungible (NFT) tokens, which widely follow the ERC-20 [1] and ERC-721 [5] standards respectively. A token contract typically must implement a set of functions and events specified in these standards. Additionally, Ether can also be called native token.
Decentralized Exchange (DEX). Similar to traditional markets, the cryptocurrency market depends on exchanges. But unlike centralized exchanges (CEX), DEX doesn’t require users to deposit digital assets into the exchange. Instead, it conducts asset transactions directly on the blockchain through smart contracts, ensuring transparency, security, and decentralization.
Flashloan. Flashloan is a unique form of cryptocurrency lending, allowing users to borrow and repay funds within a single transaction. Many DeFi protocols [7, 2, 25] currently provide flashloan function, offering everyone the opportunity to temporarily possess a large number of tokens with only a small fee. However, flashloan also facilitates numerous attacks[30, 48, 55].
Figure 2: The proportion of private adversarial transactions in 2022 and 2023.
2.3 Miner Extractable Value
Traditionally, miners determine the order of transactions in a block based on transaction fees and ending transactions in blockchain networks are transparent and publicly accessible. Therefore, users have the ability to influence the placement of their transactions by adjusting the transaction fees, allowing for front-running or back-running. The value derived from manipulating the order of transactions within a block, beyond the standard block rewards and transaction fees, is referred to as Miner Extractable Value (MEV) [34]. It has been proven to have a significant impact on the security of blockchain networks [47].
In response to MEV, researchers have proposed various countermeasures [59]. The MEV-auction solution introduced by Flashbots [20] gained popularity and aimed to facilitate efficient, decentralized, and transparent MEV extraction. Users could utilize their private service [42] to send transactions directly to miners, evading visibility within the blockchain network. However, private service providers like Flashbots on Ethereum and 48 club [14] on BSC, are also utilized by attackers to send private adversarial transactions, aiming bypass various detection effectively.
Figure 3: Kernel density estimate of interaction count for contracts created from June 2022 to June 2023
3 Data Collection
Refer to caption Contract samples Refer to captionRefer to captionExplorerRefer to captionDedaubRefer to captionClassifierRefer to captionTransaction DataRefer to captionCreator AttributeRefer to captionVerification StatusRefer to captionImplemented FunctionFunction CallContract TypeRefer to captionprocessNodeRPCAPIdecompilefeaturestrainRefer to caption
Figure 4: The workflow of our methodology for the adversarial contract detection.
To utilize supervised machine learning algorithms for binary classification of smart contracts, distinguishing whether they are adversarial or not, we first collect both adversarial contracts used in DeFi incidents and benign contracts unrelated to attacks.
3.1 Adversarial Contract
Data Sources. To comprehensively collect adversarial contracts, we gathered information on DeFi incidents from various publicly available sources, such as repositories on GitHub that document historic DeFi attacks [18, 23], attack alerts posted by security companies [51, 15, 16]. Subsequently, we undertake a meticulous manual review of these incidents to identify the adversarial contracts utilized by attackers. Additionally, We also conduct a thorough scrutiny of the transaction history of EOAs involved in these incidents, including the contract creators, callers, and associated transfer accounts, in order to identify any potentially overlooked attack incidents.
It should be noted that, because some attackers did not use private transactions to conceal their attacks, certain attacks have been executed by front-running bots. In response, we examined transactions that failed within the same block and subsequent blocks as the adversarial transaction, along with contract deployment in blocks preceding the attack block, aiming to trace back to the the initial adversarial contract deployed by the original attacker.
Scope and Exclusions. Front-running bots often deploy a generic contract that interprets the input data of transactions, akin to a virtual machine, to execute specific logic of the target transaction to frontrun. These contracts do not contain actual vulnerability exploitation logic and therefore are not within the scope of our detection efforts.
Furthermore, we observed that while attack incidents occurred on various blockchains such as Ethereum, BSC, Polygon, and Avalanche, Ethereum and BSC accounted for the vast majority of the data we collected. To streamline further work, we focused on attack incidents from Ethereum and BSC.
It is essential to mention that some attack incidents did not involve contract deployment. For instance, attacks resulting from private key leaks or attackers directly invokes the functions of the victim contracts using EOA. Clearly, these types of attacks fall outside the scope of our detection. Additionally, we only considered contracts containing core attacking logic. Other contracts, such as those used solely for concealing profits to avoid detection or fake token contract to facilitate attacks, were not included in our analysis.
Ultimately, we obtained a total of 304 contracts related to DeFi attacks on Ethereum and BSC from April 2020 to December 2023, with 115 on Ethereum and 189 on BSC, resulting in a cumulative loss estimated to exceed 100 million USD.
Private transactions analysis. By obtaining the labels of transactions related to DeFi incidents through Etherscan, we have the capability to scrutinize the utilization of private services in these attacks. Our analysis revealed that in the DeFi incidents on Ethereum, over 50% of the attacks involved the use of private transactions. Notably, within this specific subset, nearly all of the attacks employed private services when initiating adversarial transactions. Moreover, as depicted in Figure 2, the percentage of adversarial transactions sent using privacy services displayed a remarkable upward trend, surging from 33% in the first half of 2022 to a notable 71% in the latter half of 2023.
3.2 Benign Contract
To ensure that the benign contract samples are comprehensive and cover a wide range of contract types, we initially utilized Google’s BigQuery Ethereum service to query all transaction records from June 2022 to June 2023, and filtered transactions with the to field value of 0, resulting in a total of 417,300 contracts. These contracts do not represent all contracts created during this period. In other transactions, new contracts can also be created through internal transactions. However, based on our analysis of DeFi incidents, almost all attackers utilized standard transactions initiated by EOA to deploy adversarial contracts. Therefore, we exclusively focus on these contracts.
Furthermore, we conducted interaction count statistics for these contracts using the API provided by Etherscan[19]. The interaction encompass both normal transactions and internal transactions. Due to API constraints, we could retrieve a maximum of 10,000 transactions for each transaction type. The results are depicted in Figure 3. In the entirety of contracts within the specified year, the majority exhibited low interaction counts, with over a quarter of contracts having only a single transaction, which typically corresponds to the contract deployment itself, without any subsequent interactions.
To constitute our benign contract samples, we selected the top 13,000 contracts with the highest interaction counts. The minimum number of interactions among them is close to 1000. We believe that contracts with a high number of interactions are less likely to be associated with attacks, and they exhibit a certain degree of representativeness.
4 Details
In this section, we present the detailed methodology for detecting adversarial contracts, with the complete workflow depicted in Figure 4. We first introduce the six categories of features we have selected, detailing their relevance and potential impact on the detection in §4.1, then present the specific methods employed for feature extraction, focusing on implementation details in §4.2, and finally we outlines the process of building classifier using machine learning algorithms in §4.3
4.1 Feature Selection
Table 2 shows a subset of features we selected to classify contracts based on DeFi attack behaviors.
4.1.1 Creator Attribute
Creator Attribute reflects the features of the contract creator. The nonce indicates the number of transactions sent from the contract creator before the deployment of the contract. A higher nonce may suggest a more active EOA address, while attackers often use a brand-new address for their attacks. We also obtain information about the contract creator’s fund source. Typically, attackers acquire funds anonymously, such as through mixers (e.g, Tornado Cash), whereas regular users are more likely to source their funds from KYC-required centralized exchanges.
4.1.2 Transaction Data
Transaction Data includes information related to the transaction that created the contract. In cases where the contract’s constructor is payable, the sender has the capability to directly transfer native tokens, such as ETH on Ethereum, to the contract address during the contract’s creation. Typically, attackers do not need to transfer native tokens to the adversarial contract.
Within the context of the contract deployment, the input data represents the bytecode. The gas consumption depends on the specific content of the data. Consequently, the length of the input data and the corresponding gas consumption could serve as indicators of the complexity of the contract.
We do not select gas price as a feature because it fluctuates dynamically with network congestion. What may be considered a high gas price at one time may not be so at another. Additionally, attackers usually do not increase the gas price to prevent contract deployment from being frontrun.
Table 1: Flashloan Callback and Token-related Function
Flashloan Callback ERC-20 DEX
uniswapV2Call
totalSupply
uniswapV3FlashCallback
balanceOf
uniswapV3SwapCallback
transfer
swap
pancakeCall
allowance
sync
pancakeV3FlashCallback
approve
skim
DPPFlashLoanCall
transferFrom
getReserves
DVMFlashLoanCall
mint
token0
executeOperation
burn
token1
receiveFlashLoan
withdraw
onFlashLoan
deposit
BiswapCall
Table 2: Selected Features
Feature Type Name Description
Creator Attribute
Nonce
Number of transactions sent from the contract creator
Fund Source
Source of fund for the contract creator
Transaction Data
Value
Whether the native token (like ETH on Ethereum) was transferred to the contract address during contract creation
InputDataLength
Length of transaction input data
GasUsed
Amount of gas consumed by the transaction
Verification Status
Verified
Whether the contract is verified (open source)
Implemented Function
PublicFuncNumber
Number of public functions implemented
PrivateFuncNumber
Number of private functions implemented
TotalFuncNumber
Total number of functions (public and private) implemented
PublicFuncProportion
Proportion of public functions in all functions
FlashloanNumber
Number of flashloan callback functions implemented
FlashloanProportion
Proportion of flashloan callback functions in all public functions
Function Call
ExternalCallNumber
Number of external calls
InternalCallNumber
Number of internal calls
TotalCallNumber
Total number of external and internal calls
ExternalCallProportion
Proportion of external calls in all calls
DelegateCallNumber
Number of delegatecall
TokenCallNumber
Number of external calls related to tokens
TokenCallProportion
Proportion of token-related calls in all external calls
MaxTokenCall
Maximum number of token-related external calls within a single public function
AvgTokenCall
Average number of token-related external calls per public function
Contract Type
TokenContract
Whether the contract implements a specific ERC token standard, such as ERC-20 and ERC-721
ProxyContract
Whether the contract is a proxy contract
SelfDestructive
Whether the contract is self-destructive
4.1.3 Verification Status
Contracts can be verified by providing their source code and compiler settings to platforms like Etherscan. Verified contracts are open source, allowing anyone to view their source code. Many development frameworks, such as Foundry[22], offer options to verify contracts upon deployment. Generally, benign contracts tend to be verified to demonstrate their security and trustworthiness. Attackers, on the other hand, typically do not verify the adversarial contracts they deployed, exposing their malicious intent.
4.1.4 Implemented Function
This category covers features related to functions implemented in the contract. It includes the number of public functions, private functions, total functions, and the proportion of public functions. In particular, we incorporate the number of flashloan callback functions and their proportion among all public functions. This choice is motivated by the fact that many attack scenarios may require significant token holdings, and attackers may not possess these tokens, necessitating the use of flashloans to borrow tokens and execute complex attack strategies within corresponding callback functions. Additionally, since lending pool sizes may be limited, attackers might need to borrow tokens from multiple sources, resulting in multiple flashloan callback functions. The flashloan callback functions are referred to in Table 1.
4.1.5 Function Call
The Function Call provides insights into the specific calls of each function in the contract, including the number of external calls and internal calls, the total number of calls, and the ratio of external calls. It’s important to clarify that while we refer to them as internal calls, they are distinct from external calls, as they utilize the jump opcode instead of the call opcode. We also include the number of delegatecall.
In addition, we consider external calls related to ERC-20 tokens, including their count and proportion. The specific functions are outlined in Table 1. Generally, the objective of DeFi attack is to gain economic benefits, i.e., acquire valuable tokens and eventually convert them into native tokens for withdrawal. Therefore, external calls within adversarial contracts often involve token-related functions.
While the withdraw and deposit functions are not ERC-20 standard functions, they are implemented for wrapped native tokens. Among the functions within DEX implementations, In addition to swap related functions, we also include some UniswapV2 protocol functions, such as sync and skim, as they are also common functions used in some attacks.
Furthermore, We introduce two additional features, MaxTokenCall and AvgTokenCall. MaxTokenCall represents the maximum number of token-related calls within a single public function execution flow, which helps identify functions with exceptionally high token interactions. AvgTokenCall, on the other hand, indicates the average number of token-related calls per public function execution flow, providing a measure of the average token interaction level.
4.1.6 Contract Type
Token contracts and proxy contracts are widely prevalent on the blockchain, while adversarial contracts typically do not fall into these categories. Additionally, some attackers have a tendency to self-destruct their contracts subsequent to the completion of the attack. As a result, we include the contract type as one of the distinctive features.
4.2 Feature Extraction
4.2.1 Transaction Data Retrieval
In the first step of feature extraction, we utilize an RPC (Remote Procedure Call) provider to send direct requests to nodes within the blockchain. This allows us to obtain the raw data of transactions corresponding to the contract deployment, thereby acquiring essential features such as nonce, value, input data length, and gas consumption. Additionally, the transaction sender, input data and created contract address can also be obtained, which will be utilized in subsequent phases of feature extraction.
4.2.2 Explorer Data Retrieval
Subsequently, we leverage the APIs provided by blockchain explorers, namely Etherscan and Bscscan, to easily retrieve information about the contract’s verification status.
We employ Algorithm 1 to obtain the labels of the fund source. We begin by retrieving the complete transaction history of the contract creator, encompassing both normal and internal transactions. We then identify the earliest transaction where native tokens were transferred to the contract creator’s address, considering its sender as the contract creator’s fund source.
The obtained address at this stage cannot serve as a feature directly. Therefore, we combine the address with the publicly available address labels and tags from Etherscan and Bscscan. For each fund source address corresponding to our contract samples, we construct a mapping table associating addresses with their respective labels. This mapping table enables us to extract explicit fund source features for all fund source addresses.
We categorize fund source addresses into five types:
  1. 1)
    Safe: Denotes fund sources with relatively secure origins, such as KYC-required centralized exchanges, which are less likely to be used by attackers.
  2. 2)
    Anonymous: Signifies fund sources with anonymous origins, including mixers and some KYC-free exchanges, which are more susceptible to be used for malicious purposes.
  3. 3)
    Bridge: Indicates funds originating from cross-chain transfers. We cannot definitively determine if their source is anonymous or not, so they are considered potentially usable for attacks.
  4. 4)
    Unknown: Addresses lacking specific labels but possessing a substantial history of transfer transactions. We cannot categorize them accurately.
  5. 5)
    Transfer: Addresses with minimal transaction activity, suggesting that the fund source is an ordinary address used for basic transfer.
Input: The contract creator $c$
Output: The fund source label of $c$
1 init
2       $C_{\textnormal{hop}}\leftarrow$ 2
3       $C_{\textnormal{threshold}}\leftarrow$ 10000
4       $D_{\textnormal{label}}\leftarrow$label1: [address1, address2], label2: [address3, address4], …}
5
6
7while $C_{\textnormal{hop}}$geq 0 do
8       $T_{\textnormal{xn}}\leftarrow$ First normal transaction of $c$ with a native token transfer in.
9       $T_{\textnormal{xi}}\leftarrow$ First internal transaction of $c$ with a native token transfer in.
10       $T_{\textnormal{x}}\leftarrow$ The transaction with the smaller blocknum between $T_{\textnormal{xn}}$ and $T_{\textnormal{xi}}$.
11       $f_{\textnormal{address}}\leftarrow$ The sender address of $T_{\textnormal{x}}$.
12       if $f_{\textnormal{address}}$ in $D_{\textnormal{label}}$ then
13             return The label of $f_{\textnormal{address}}$ in $D_{\textnormal{label}}$
14       end if
15      $C_{\textnormal{hop}}$ = $C_{\textnormal{hop}}$ 1;
16       if $C_{\textnormal{hop}}$ ==absent$=$ = 0 then
17             $n\leftarrow$ Historical transaction count of $f_{\textnormal{address}}$
18             if $n$ < $C_{\textnormal{threshold}}$ then
19                   return’Transfer’
20             end if
21            return’Unknown’
22       end if
23      $c\leftarrow f_{\textnormal{address}}$
24 end while
Algorithm 1 Fund Source Labeling Algorithm
To further refine the accuracy of fund source feature labels, in the case where we identify the fund source address as belonging to the Transfer category, we will continue to trace it once, that is, conduct an additional lookup to ascertain the label of the fund source corresponding to the fund source address again. This step is crucial to prevent attackers from using intermediary address transfers to evade detection of their malicious fund source, thereby ensuring more precise identification of the fund source.
4.2.3 Decompilation
Feature Model Accuracy Precision Recall F1-Score FPR
Transaction features Logistic Regression 0.9804 0.5053 0.9057 0.6486 0.0181
Decision Trees 0.9845 0.6667 0.4528 0.5393 0.0046
Random Forest 0.9883 0.7500 0.6226 0.6804 0.0042
MLP 0.9804 0.5059 0.8113 0.6232 0.0162
LightGBM 0.9872 0.7111 0.6038 0.6531 0.0050
Contract features Logistic Regression 0.9453 0.2629 0.9623 0.4130 0.0550
Decision Trees 0.9879 0.6721 0.7736 0.7193 0.0077
Random Forest 0.9917 0.7925 0.7925 0.7925 0.0042
MLP 0.9876 0.6429 0.8491 0.7317 0.0096
LightGBM 0.9925 0.8367 0.7736 0.8039 0.0031
Hybrid features Logistic Regression 0.9872 0.6234 0.9057 0.7385 0.0112
Decision Trees 0.9925 0.8511 0.7547 0.8000 0.0027
Random Forest 0.9970 0.9245 0.9245 0.9245 0.0015
MLP 0.9951 0.8448 0.9245 0.8829 0.0035
LightGBM 0.9977 0.9434 0.9434 0.9434 0.0012
Table 3: Performance Metrics for Different Features and Models
Finally, we utilize tools provided by Dedaub[17] to decompile the contract bytecode, allowing us to analyze and extract the remaining features. Regardless of whether the contract is verified, to facilitate subsequent unified processing, we will perform the decompilation.
Specifically, we first decompile the contract’s bytecode (transaction input data) and runtime code (deployed on the chain) to obtain pseudocode. Then, we use regular expressions to extract information about implemented functions, including their names and visibility, thereby obtaining all features for the Implemented Function category. Subsequently, we also utilize regular expressions to further analyze the function calls within each function, obtaining features for the Function Call category, and determine whether the contract can self-destruct.
It’s worth noting that Dedaub’s decompilation attempts to automatically resolve function signatures, but the results may not always be complete. Therefore, we use both function names and function selectors to detect the implementation of flashloan callback functions and token-related calls.
To extract the MaxTokenCall and AvgTokenCall features, we analyze the execution flow of each public function, which means detecting the internal calls and potential flashloan function calls, and then counting token related calls involved in them.
Furthermore, to determine whether the contract is a token contract, we inspect the implemented function to check if it fully implements all functions of a specific token ERC standard. Specifically, we implement detection for ERC-20[1], ERC-721[5], and ERC-1155[4] token contract standards. (Other token ERC standards ERC-777[3] and ERC-4626[9] are backward-compatible with ERC-20)
For identifying proxy contracts, despite various proxy patterns, most proxy contracts, following ERC-1967[6], typically saves the logic contract’s address in a specific storage slot. Therefore, if a contract involves a delegatecall and the contract creation code contains the slot value, we consider it a proxy contract, which covers the majority of cases.
4.3 Classifier Building
Once all the required features have been collected, we proceed to train a supervised learning classifier using the previously gathered datasets of benign and adversarial contracts.
Due to the limited quantity of adversarial samples, to mitigate the issue of class imbalance that can bias model predictions, we employed an oversampling technique known as ADASYN (Adaptive Synthetic Sampling). This approach synthesizes additional adversarial samples to bring the number of attack samples in the training set closer to that of benign samples.
For the fund source feature, which consists of categorical data with low cardinality, we applied one-hot encoding. Additionally, to ensure that features of different dimensions are more comparable in numerical terms and to enhance the classifier’s accuracy, we standardized the data using Z-score normalization. This process centers each feature’s values around 0 with a standard deviation of 1.
Finally, when selecting an appropriate supervised machine learning algorithm to construct the classifier, we explore various methods. These methods include classical approaches such as Logistic Regression and Decision Trees. We also try ensemble learning approaches, such as Random Forest based on bagging and the LightGBM algorithm based on boosting. In addition, we considered the use of a classic neural network architecture, the Multilayer Perceptron (MLP), as one of the candidate models.
5 Evaluation
5.1 Metrics
In addition to commonly used binary classification performance metrics such as accuracy, precision, recall, F1-Score, we also calculate the False Positive Rate (FPR) metric, which holds particular significance in our scenario. A high FPR implies that the classifier incorrectly identifies benign contracts as adversarial, leading to substantial manual audit costs.
A distinguishing aspect of our research compared to traditional supervised machine learning problems is the explicit chronological relationship between contract deployment. Consequently, traditional cross-validation techniques may not be suitable for our dataset, as they assume that each data is independent. And our objective is to train a model based on existing data to detect potential future attacks, so for sustainability reasons, we order the contract samples chronologically, allocating the first 80% for training and the remaining portion for testing to calculate relevant metrics, instead of relying on cross-validation.
5.2 Performance
Table 3 presents our test results. Overall, the LightGBM model outperforms others, demonstrating the best performance across all metrics when using all features. It exhibits a remarkably high detection rate for adversarial samples, with an F1-Score of 0.9434, while maintaining an exceptionally low FPR of 0.12%. Other models also perform well, indicating the applicability of our approach to the dataset.
To the best of our knowledge, this is the first academic work on the detection of adversarial contracts utilized in DeFi attack. The only similar research made public so far is by the Forta Network[21], which also uses machine learning. However, their approach differs as it focuses solely on low-level bytecode. We used their method as a baseline for comparison.
Model Precision Recall F1-Score
forta 0.8778 0.5520 0.6208
ours 0.9434 0.9434 0.9434
Table 4: Model Performance Metrics
As shown in Table 4, our model outperforms Forta’s most recently published model (released on February 6, 2023) across all performance metrics. This highlights the superiority of our method.
5.3 Feature Importance
To assess the impact of different feature categories, we divided the features into two major classes: one focused on transaction-related attributes (Creator Attribute, Transaction Data, Verification Status), and the other related to the contract’s code itself (Implemented Function, Function Call, Contract Type). We conducted tests using only one feature category at a time, and the results are presented in Table 3.
Clearly, when compared to using a combination of features, training classifiers with a single category of features leads to a certain degree of performance degradation for all models. This indicates that each category of features plays a valuable role in the detection of adversarial contracts. Additionally, Figure 5 illustrates the feature importance of all features for the LightGBM model.
Figure 5: Top 15 LightGBM Feature Importance
6 Limitations and Future Work
In this section, we discuss the limitation as well as the possible future work.
Detection Scope. Our current method has certain limitations in the scope of detection. In terms of feature selection, our emphasis is DeFi attacks involving financial losses, specifically tailored to common attack scenarios. Consequently, we may miss attacks that do not result in financial losses and certain specialized attack types. For instance, during evaluation, we found that a small number of undetected attacks related to NFTs, for which we did not have specifically designed features. Moreover, we have observed that some DeFi attackers, similar to front-running bots, encapsulate function calls related to attacks within calldata. As the attacking logic employing this approach is not included in the contract, it also poses difficulties for detection.
Performance. The current feature extraction process relies on RPC providers and blockchain explorer APIs. While these components can be replaced by running our custom nodes, there is currently a lack of robust open-source decompilation tools. Solely relying on Dedaub’s online decompilation service results in slow response times, making real-time detection impractical.
In the future, we consider adjusting certain features to accommodate evolving attack types. Furthermore, there is room for constructing our own decompilation module. This would not only improve feature extraction performance but also allow for the extraction of more comprehensive features from bytecode.
7 Related Work
In this section, we discuss existing works that are related to our work. In particular, we first discuss existing works for adversarial transaction detection, then present the papers related to abnormal contract detection.
7.1 Adversarial Transaction Detection
Some researches[38, 50, 31, 57] modify EVM to obtain more detailed transaction execution information and provide scalable interfaces for online detection of various types of adversarial transactions. Sereum[50] uses taint analysis to monitor runtime data flows during the execution of transactions to detect reentrancy attacks. EthScope[57] is a scalable Ethereum attack detection framework which stores historical data into an cluster database and adds dynamic taint analysis to the client to analyze transactions.
On the other hand, some researches[60, 63, 54, 58, 46] achieve adversarial transaction detection through forensic analysis without modifying EVM. TxSpector[60] is a logic-driven framework designed to analyze Ethereum transactions, identifying attacks according to user-defined detection rules by encoding the transaction trace into logical relationships. Qin et al.[46] propose an Execution Property Graph and a graph traversal technique which can swiftly identify potential adversarial transactions efficiently.
7.2 Abnormal Contract Detection
Forta‘s [21] method combines Natural Language Processing (NLP) techniques to identify adversarial contracts at the bytecode level, and apart from it, there is no existing work that specifically focuses on the detection of adversarial contracts. The majority of abnormal contract detection focuses on Ponzi schemes and honeypots. They share similarities in methods and can be categorized into code similarity detection[27, 52], machine learning[61, 40, 41, 39, 29, 33], and symbolic execution techniques[32, 53].
Ponzi scheme. Ponzi scheme is a financial fraud that attracts users with promises of high profits. With the emergence of blockchain technology, it has provided many opportunities for scammers. Zheng et al.[32] constructed a larger dataset of Ponzi scheme than previous work[33] and proposed a Multi-view Cascade Ensemble method for detecting smart Ponzi schemes on Ethereum.
Honeypots. Honeypots are a distinct type of contracts designed to entice victims into traps, employing seemingly vulnerable structures that conceal hidden risks rather than relying on traditional deceptive methods. Torres et al.[53] developed a taxonomy of honeypot techniques and built a tool called HONEYBAGER that employs symbolic execution and well defined heuristics to expose honeypots.
8 Conclusion
In this paper, we propose a new direction for detecting DeFi attacks, i.e., detecting adversarial contracts instead of adversarial transactions. We build the first dataset for adversarial contracts. Based on DeFi attack patterns, we build a machine learning classifier that can effectively distinguish adversarial contracts from benign ones. The evaluation show that our method performs exceptionally well. The F1-Score for LightGBM-based classifier is 0.9434, with a remarkably low false positive rate of only 0.12