### 02 Feb 2024 v2

# LookAhead: Preventing DeFi Attacks via Unveiling Adversarial Contracts

Abstract.
DeFi incidents stemming from various smart contract vulnerabilities have culminated in financial damages exceeding 3 billion USD. The attacks causing such incidents commonly commence with the deployment of adversarial contracts, subsequently leveraging these contracts to execute adversarial transactions that exploit vulnerabilities in victim contracts. Existing defense mechanisms leverage heuristic or machine learning algorithms to detect adversarial transactions, but they face significant challenges in detecting private adversarial transactions. Namely, attackers can send adversarial transactions directly to miners, evading visibility within the blockchain network and effectively bypassing the detection.
In this paper, we propose a new direction for detecting DeFi attacks, i.e., detecting adversarial contracts instead of adversarial transactions, allowing us to proactively identify potential attack intentions, even if they employ private adversarial transactions. Specifically, we observe that most adversarial contracts follow a similar pattern, e.g., anonymous fund source, closed-source, frequent token-related function calls. Based on this observation, we build a machine learning classifier that can effectively distinguish adversarial contracts from benign ones.
We build a dataset consists of features extracted from 269 adversarial contracts and 13,000 benign contracts. Based on this dataset, we evaluate different classifiers, the results of which show that our method for identifying DeFi adversarial contracts performs exceptionally well. For example, the F1-Score for LightGBM-based classifier is 0.9541, with a remarkably low false positive rate of only 0.15%.
copyright: noneconference: ; October 14–18, 2024; Salt Lake City, USA
1. Introduction
Decentralized Finance (DeFi) has gained significant prominence and is evolving rapidly within the blockchain ecosystem in recent years, giving rise to a diverse array of DeFi applications such as decentralized exchanges (DEXs) (UniswapV3, 3, 44), lending platforms (AAVE, 2, 52), and derivatives (dydx, 25). The total value locked (TVL) within these DeFi protocols exceeded \$50 billion in early 2024 (DefiLlama, 62). Among various blockchain networks, Ethereum (wood2014ethereum, 68) and Ethereum Virtual Machine (EVM)-compatible chains (Arbitrum, 4, 10, 45, 47) have a thriving DeFi ecosystem, playing a pivotal role in the landscape.
Refer to captionAttackerRefer to caption Victim contract(s) 1attack2profit
(a) Attacking directly.
Refer to captionAttackerRefer to caption Victim contract(s) Refer to captionAdversarial contract1deploy2callattack3profit4
(b) Attacking through adversarial contracts.
Figure 1. Two common patterns for DeFi attacks: (a) directly interacts with the victim contracts (b) interacts through adversarial contracts.
At the heart of DeFi are smart contracts, which are self-executing programs deployed on blockchains, enabling trustless and transparent financial interactions without the need for intermediaries. They could be invoked by users through transactions and autonomously execute predefined actions. However, like traditional software, they are fundamentally code-based programs, making them susceptible to potential security risks. Indeed, DeFi incidents resulting from various smart contract vulnerabilities have caused over 3 billion USD in losses (DeFiHackLabs, 24, 42, 75). Figure 1 depicts two common attack patterns:
  1. 1)
    Attacking directly. The attacker directly invokes the functions of the victim contracts (TreasureDAOHack, 6, 59). However, in the presence of complex attacking logic involving multiple transactions, there is a potential risk of half-way failure. This situation arises when part of the attack is executed but the rest cannot proceed (e.g., due to transaction conflictions).
  2. 2)
    Attacking through adversarial contracts. The attacker first deploys an adversarial contract that contains specific attack scripts for the victim contracts, and then initiates the attack by calling the entry function of the adversarial contract (BeanstalkHack, 5, 11, 46, 58). This pattern allows the attacker to execute complex attacking logic in a single transaction, hence avoid half-way failure due to the atomic nature of a transaction. Most DeFi attacks follow this pattern.
Existing defense mechanisms leverage heuristic (qin2023towards, 49, 70, 72, 76) or machine learning (gai2023blockchain, 37, 64) algorithms to detect adversarial transactions, but they face significant challenges in detecting private adversarial transactions. Namely, attackers can send adversarial transactions directly to miners, evading visibility within the blockchain network and effectively bypassing the detection. Based on our analysis of historical DeFi incidents, we found that the proportion of adversarial transactions sent using privacy services has exhibited a significant upward trend, rising from 33% in the first half of 2022 to a notable 63% in the latter half of 2023.
In this paper, we propose a new direction for detecting DeFi attacks, i.e., detecting adversarial contracts instead of adversarial transactions, allowing us to proactively identify potential attack intentions, even if they employ private adversarial transactions.
We first collected 269 adversarial contracts used in DeFi incidents between April 2020 and December 2023 from both the Ethereum and Binance Smart Chain (BSC). We also collected the top 13,000 contracts on Ethereum from June 2022 to June 2023, to serve as our benign contract samples. These data are made publicly available to facilitate future research in this domain.
We observe that most adversarial contracts follow a similar pattern, e.g., anonymous fund source, closed-source, frequent token-related function calls. Based on this observation, we extract six categories of features from each contract, such as attributes of contract creator, implemented functions in contract, and contract types. We then build classifiers using multiple supervised machine learning algorithms and evaluate them comprehensively. The results show that our method for identifying DeFi adversarial contracts performs exceptionally well. For example, the F1-Score for LightGBM-based classifier is 0.9541, with a remarkably low false positive rate of only 0.15%.
We summarize our contributions as follows:
  • A new direction for detecting DeFi attacks, i.e., detecting adversarial contracts instead of adversarial transactions.
  • The first dataset for adversarial contracts.
  • A comprehensive evaluation of different classifiers.
2. Background
In this section, we outline the required background to better understand our work.
2.1. Ethereum
Blockchain is a distributed ledger built on a peer-to-peer (P2P) network, maintained through specific consensus mechanisms to ensure the consistency of the ledger. Ethereum (wood2014ethereum, 68) is one of the most renowned blockchain platforms. It introduces Ethereum Virtual Machine (EVM) to support smart contracts and has spawned a series of EVM-compatible chains.
Account. Accounts are entities that can hold Ether and send transactions. Ethereum has two types of accounts: External Owned Accounts (EOAs) and Contract Accounts (CAs). EOAs are controlled by users’ private keys, while CAs contain corresponding code and storage. Both types of accounts have their address, balance, and nonce. The address of an EOA is derived from the hash of the public key, whereas the address of a CA is usually related to its creator.
Transcation. Ethereum is often regarded as a transaction-driven state machine, where transactions are initiated by EOA, signed, and ultimately update the Ethereum network’s state. After a transaction is sent, it’s first placed in the mempool, a transparent and publicly accessible pool filled with pending transactions, waiting for the block producer to select and finalize them. Transactions are typically used for transferring Ether, invoking contracts, or creating contracts. In the case of contract deployment, the to field of the transaction is 0. During the execution of a normal transaction, interacted contracts may trigger additional message calls, commonly known as internal transactions.
Smart Contract. From a high-level perspective, smart contracts can be viewed as instantiated objects stored on the blockchain and executed within the EVM. Due to the nature of blockchain, once deployed, the code within a smart contract cannot be changed. Smart contracts are typically written in high-level programming languages such as Solidity (dannen2017introducing, 22). The source code is compiled to bytecode, which can be executed within the EVM. The EVM is a stack-based virtual machine that supports a set of Turing-complete instructions. Bytecode mainly consists of two parts, creation code and runtime code. Creation code is used for initialization (not deployed on the blockchain) and is executed only once. Runtime code is the primary code segment executed when the contract receives a message call.
2.2. DeFi
DeFi is a financial ecosystem that operates on a blockchain network. It is built upon smart contracts and aims to provide traditional financial services in a more open and transparent manner. DeFi has also introduced some unique financial concepts, such as tokens, decentralized exchange, and flashloan.
The tokens. There are numerous cryptocurrencies created by smart contracts on Ethereum, known as tokens. These tokens serve as the cornerstone of the DeFi ecosystem, facilitating the diversified development of applications and services within DeFi. Tokens come in two main types: fungible tokens and non-fungible tokens (NFT), which widely follow the ERC-20 (ERC20, 28) and ERC-721 (ERC721, 30) standards respectively. A token contract typically must implement a set of functions and events specified in these standards. Furthermore, Ether, which is used to reward contribution in the consensus, is also referred to as the native token.
Decentralized Exchange (DEX). Similar to traditional markets, the cryptocurrency market depends on exchanges. But unlike centralized exchanges (CEX), DEX doesn’t require users to deposit digital assets into the exchange. Instead, it conducts asset transactions directly on the blockchain through smart contracts, ensuring transparency, security, and decentralization. The most prevalent DEXs (UniswapV3, 3, 44) utilize the Automated Market Maker (AMM) model, which relies on mathematical formulas to dynamically adjust the prices of digital assets, diverging from the traditional order book approach used in CEXs. In this model, the core contracts known as liquidity pools are funded by entities called liquidity providers, allowing users to continuously exchange assets.
Flashloan. Flashloan is a unique form of cryptocurrency lending that leverages the atomic nature of the transaction, where a transaction is either fully executed or not executed at all, allowing users to borrow and repay cryptocurrency in one transaction. To utilize the flashloan, specific callback functions must be implemented in the borrowing smart contracts. Many DeFi protocols (AAVE, 2, 3, 25) currently provide flashloan function, offering everyone the opportunity to temporarily possess a large number of tokens with only a small fee. However, flashloan also facilitates numerous attacks (cao2021flashot, 13, 51, 65).
Figure 2. The proportion of private adversarial transactions in 2022 and 2023.
2.3. Miner Extractable Value
Traditionally, miner determine the order of transactions in a block based on transaction fees. Since the mempool is publicly accessible, users have the ability to influence the placement of their transactions by adjusting the transaction fees, allowing for front-running (eskandari2020sok, 32, 60) or back-running. The value derived from manipulating the order of transactions within a block, beyond the standard block rewards and transaction fees, is referred to as Miner Extractable Value (MEV) (daian2020flash, 21). It has been proven to have a significant impact on the security of blockchain networks (qin2022quantifying, 50).
Front-running. Front-running is a fundamental means used to extract MEV, typically employed in arbitrage, liquidation, and sandwich attack. Front-runners usually deploy a contract on the blockchain, akin to a virtual machine, capable of interpreting the input data of transactions. When the front-running bot detects a victim transaction that meets specific conditions, it constructs input data to execute the specific logic for profit. On the other hand, front-running has also been considered as a means in the intrusion prevention systems (qin2023blockchain, 48, 73), attempting to frontrun adversarial transactions to prevent the attacks from occurring.
Private Service. In response to MEV, researchers have proposed various countermeasures (yang2022sok, 71). The MEV-auction solution introduced by Flashbots (Flashbots, 34) gained popularity (weintraub2022flash, 67). Users could utilize their service (lyu2022empirical, 43) to send private transactions directly to miners, evading exposure in the mempool. However, private service providers like Flashbots on Ethereum and 48 club (bnb48, 1) on BSC, are also utilized by attackers to send private adversarial transactions, aiming bypass various detection effectively.
Figure 3. Kernel density estimate of interaction count for contracts created from June 2022 to June 2023.
3. Data Collection
Refer to caption Contract samples Refer to captionRefer to captionExplorerRefer to captionDedaubRefer to captionClassifierRefer to captionTransaction DataRefer to captionCreator AttributeRefer to captionVerification StatusRefer to captionImplemented FunctionFunction CallContract TypeRefer to captionprocessNodeRPCAPIdecompilefeaturestrainRefer to caption
Figure 4. The workflow of our methodology for the adversarial contract detection.
Category Count (Proportion)
Total Attacks 109
Involving Private 54 (49.54%)
Adversarial Tx Private 52 (96.30% of Private)
Table 1. Analysis of private transactions in the DeFi incidents on Ethereum
To utilize supervised machine learning algorithms for binary classification of smart contracts, distinguishing whether they are adversarial or not, we first collect both adversarial contracts used in DeFi incidents and benign contracts unrelated to attacks.
3.1. Adversarial Contract
Data Sources. To comprehensively collect adversarial contracts, we gathered information on DeFi incidents from various publicly available sources, such as repositories on GitHub that document historic DeFi attacks (DeFiHackLabs, 24, 42), attack alerts posted by security companies (blocksec, 8, 15, 55). Subsequently, we undertake a meticulous manual review of these incidents to identify the adversarial contracts utilized by attackers. Additionally, We also conduct a thorough scrutiny of the transaction history of EOAs involved in these incidents, including the contract creators, callers, and associated transfer accounts, in order to identify any potentially overlooked attack incidents.
It should be noted that, because some attackers did not use private transactions to conceal their attacks, certain attacks have been executed by front-running bots. In response, we examined transactions that failed within the same block and subsequent blocks as the adversarial transaction, along with contract deployment in blocks preceding the attack block, aiming to trace back to the the initial adversarial contract deployed by the original attacker.
Scope and Exclusions. Due to the contracts deployed by front-running bots do not contain actual vulnerability exploitation logic and therefore are not within the scope of our detection efforts. Furthermore, we observed that while attack incidents occurred on various blockchains, Ethereum and BSC accounted for the vast majority of the data we collected. To streamline further work, we focused on attack incidents from Ethereum and BSC. Additionally, we only considered contracts containing the core attacking logic. Other contracts, such as those used solely for concealing profits to avoid detection or fake token contract to facilitate attacks, were not included in our analysis.
Ultimately, we obtained a total of 269 contracts related to DeFi attacks on Ethereum and BSC from April 2020 to December 2023, with 109 on Ethereum and 160 on BSC, resulting in a cumulative loss estimated to exceed 1 billion USD.
Private transactions analysis. Transactions with labels of MEV Transactions or Flashbots in Etherscan indicate that they bypassed the Ethereum mempool and were sent directly to the block producer. By obtaining the labels of transactions related to DeFi incidents through Etherscan, we have the capability to scrutinize the utilization of private transactions in these attacks. As shown in table 1, our analysis revealed that in the DeFi incidents on Ethereum, about 50% of the attacks involved the use of private transactions. Notably, within this specific subset, nearly all of the attacks employed private services when initiating adversarial transactions. Moreover, as depicted in Figure 2, the percentage of adversarial transactions sent using privacy services displayed a remarkable upward trend, surging from 33% in the first half of 2022 to a notable 63% in the latter half of 2023.
3.2. Benign Contract
To ensure that the benign contract samples are comprehensive and cover a wide range of contract types, we initially utilized Google’s BigQuery Ethereum service to query all transaction records from June 2022 to June 2023, and filtered transactions with the to field value of 0, resulting in a total of 417,300 contracts. These contracts do not represent all contracts created during this period. In other transactions, new contracts can also be created through internal transactions. However, based on our analysis of DeFi incidents, almost all attackers utilized standard transactions initiated by EOA to deploy adversarial contracts. Therefore, we exclusively focus on these contracts.
Furthermore, we conducted interaction count statistics for these contracts using the API provided by Etherscan (etherscan, 33). The interaction encompass both normal transactions and internal transactions. Due to API constraints, we could retrieve a maximum of 10,000 transactions for each transaction type. The results are depicted in Figure 3. In the entirety of contracts within the specified year, the majority exhibited low interaction counts, with over a quarter of contracts having only a single transaction, which typically corresponds to the contract deployment itself, without any subsequent interactions.
To constitute our benign contract samples, we selected the top 13,000 contracts with the highest interaction counts. The minimum number of interactions among them is close to 1,000. We believe that contracts with a high number of interactions are less likely to be associated with attacks, and they exhibit a certain degree of representativeness.
4. Details
In this section, we present the detailed methodology for detecting adversarial contracts, with the complete workflow depicted in Figure 4. We first introduce the six categories of features we have selected, detailing their relevance and potential impact on the detection in §4.1, then present the specific methods employed for feature extraction, focusing on implementation details in §4.2, and finally we outlines the process of building classifier using machine learning algorithms in §4.3
4.1. Feature Selection
Feature Category Name Description
Creator Attribute
Nonce
Number of transactions sent from the contract creator
Fund Source
Source of fund for the contract creator
Transaction Data
Value
Whether the native token (like ETH on Ethereum) was transferred to the contract address during contract creation
InputDataLength
Length of transaction input data
GasUsed
Amount of gas consumed by the transaction
Verification Status
Verified
Whether the contract is verified (open source)
Implemented Function
PublicFuncNumber
Number of public functions implemented
PrivateFuncNumber
Number of private functions implemented
TotalFuncNumber
Total number of functions (public and private) implemented
PublicFuncProportion
Proportion of public functions in all functions
FlashloanNumber
Number of flashloan callback functions implemented
FlashloanProportion
Proportion of flashloan callback functions in all public functions
Function Call
ExternalCallNumber
Number of external calls
InternalCallNumber
Number of internal calls
TotalCallNumber
Total number of external and internal calls
ExternalCallProportion
Proportion of external calls in all calls
DelegateCallNumber
Number of delegatecall
TokenCallNumber
Number of external calls related to tokens
TokenCallProportion
Proportion of token-related calls in all external calls
MaxTokenCall
Maximum number of token-related external calls within a single public function
AvgTokenCall
Average number of token-related external calls per public function
Contract Type
TokenContract
Whether the contract implements a specific ERC token standard, such as ERC-20 and ERC-721
ProxyContract
Whether the contract is a proxy contract
SelfDestructive
Whether the contract is self-destructive
Table 2. Selected Features
Table 2 shows the core subset of features we selected to classify contracts based on DeFi attack behaviors, divided into six categories.
4.1.1. Creator Attribute
Figure 5. Distribution of input data length and gas consumption for both adversarial and benign contract samples.
This category encompasses features related to the contract creator. The nonce indicates the number of transactions sent from the contract creator before the deployment of the contract. A higher nonce may suggest a more active EOA address. We also obtain information about the contract creator’s fund source. To conceal their identity, attackers typically transfer initial funds from a mixer (e.g, Tornado Cash (wang2023zero, 66)) or instant exchanges (ChangeNOW, 16, 54) without the need for sign-up or KYC (Know Your Customer), to a new account for subsequent transactions. In contrast, regular users are more likely to source their funds from KYC-required centralized exchanges.
4.1.2. Transaction Data
This category pertains to information concerning the transaction that created the contract. In cases where the contract’s constructor is payable, the creator possesses the ability to directly transfer native tokens, such as ETH on Ethereum, to the contract address during the contract’s creation. The specific amount transferred is denoted by the value field within the transaction.
In the context of the contract deployment, the input data field of a transaction represents the bytecode of the contract. The gas consumption for a transaction is contingent upon the specific content of the bytecode. Consequently, the length of the input data and the corresponding gas consumption could serve as indicators of the complexity of the contract. As depicted in Figure 5, within our contract samples, it is observed that in the case of adversarial contracts, the deployment transactions typically involve less input data, with gas consumption concentrated in the range of 1.0e6 to 1.5e6 wei. Conversely, benign contracts exhibit slightly more input data, accompanied by a more widespread gas consumption, exceeding 2.0e6 in general.
We do not select gas price as a feature because it fluctuates dynamically with network congestion. What may be considered a high gas price at one time may not be so at another. Additionally, attackers usually do not need to increase the gas price to prevent contract deployment from being frontrun.
4.1.3. Verification Status
Verified Unverified
Benign 11594 (89.2%) 1406 (10.8%)
Adversarial 4 (1.5%) 265 (98.5%)
Table 3. Verification status for both benign and adversarial contract samples.
Contracts can be verified by providing their source code and compiler settings to platforms like Etherscan. Verified contracts are open source, allowing anyone to view their source code. Many development frameworks, such as Foundry (foundry, 36), offer options to verify contracts upon deployment. Generally, benign contracts tend to be verified to demonstrate their security and trustworthiness. Attackers, on the other hand, typically do not verify the adversarial contracts they deployed, exposing their malicious intent. This can also be corroborated by our contract samples presented in Table 3.
4.1.4. Implemented Function
Figure 6. Distribution of public function number and private function number for both adversarial and benign contract samples.
This category covers features related to implemented functions in the contract, such as the number of public functions, private functions, total functions, and the proportion of public functions. Upon analyzing our contract samples, as illustrated in Figure 6, it becomes apparent that benign contracts frequently incorporate more than 20 public functions, while adversarial contracts tend to implement fewer public functions. Furthermore, adversarial contracts tend to implement fewer private functions.
Figure 7. Distribution of implemented flashloan callback function number for both adversarial and benign contract samples.
In particular, we incorporate the number of flashloan callback functions and their proportion among all public functions. The flashloan callback functions are referred to in Table 4. This choice is motivated by the fact that many attack scenarios may require significant token holdings, and attackers may not possess these tokens, necessitating the use of flashloans to borrow tokens and execute complex attack strategies within corresponding callback functions. Additionally, since lending pool sizes may be limited, attackers might need to borrow tokens from multiple sources, resulting in multiple flashloan callback functions. In our contract samples, as shown in Figure 7, it is noticeable that approximately two-thirds of adversarial contracts implement flashloan callback functions, while almost all benign contracts do not involve such callbacks.
Flashloan Callback ERC-20 DEX
uniswapV2Call
totalSupply
uniswapV3FlashCallback
balanceOf
uniswapV3SwapCallback
transfer
swap
pancakeCall
allowance
sync
pancakeV3FlashCallback
approve
skim
DPPFlashLoanCall
transferFrom
getReserves
DVMFlashLoanCall
mint
token0
executeOperation
burn
token1
receiveFlashLoan
withdraw
onFlashLoan
deposit
BiswapCall
Table 4. Flashloan Callback and Token-related Function
4.1.5. Function Call
This category provides insights into the specific calls of each function in the contract, including the number of external and internal calls, the total number of calls, the ratio of external calls, and the number of delegatecall. It’s important to clarify that while we refer to them as internal calls, they are distinct from external calls because they utilize the jump opcode instead of the call opcode, signifying a fundamental difference in the underlying EVM execution. In our contract samples, as shown in Figure 8, over half of adversarial contracts involve more than 20 external calls, while most benign contracts have fewer than 10 external calls. Additionally, benign contracts tend to have more internal calls compared to adversarial contracts.
We have also taken into consideration external calls related to ERC-20 tokens, including their count and proportion. The specific functions are outlined in Table 4. Generally, the objective of DeFi attack is to gain financial benefits, i.e., acquire valuable tokens and eventually convert them into native tokens for withdrawal. Therefore, external calls within adversarial contracts often involve token-related functions.
While the withdraw and deposit functions are not ERC-20 standard functions, they are implemented for wrapped native tokens. Among the functions within DEX implementations, in addition to swap related functions, we also include some UniswapV2 (UniswapV2, 63) protocol functions, such as sync and skim, as they are also common functions used in some attacks. In our sample contracts, as shown in Figure 9, it is evident that over 70% of benign contracts do not have token-related calls. In contrast, almost all adversarial contracts involve token-related calls, with over half of them having 16 or more token-related calls.
Furthermore, we introduce two novel features to assess function interactions with tokens: MaxTokenCall and AvgTokenCall. To better understand their significance, it’s essential to first consider the concept of the function execution flow. The flow encompasses the complete process a function undergoes during its operation. Specifically, when a public function is executed, it often involves invoking other implemented functions. MaxTokenCall is designed to capture the peak level of token-related interactions within a single public function execution flow, which helps identify functions with exceptionally high token interactions. On the other hand, AvgTokenCall indicates the average number of token-related calls per public function execution flow, providing a measure of the average token interaction level.
Figure 8. Distribution of external and internal call number for both adversarial and benign contract samples.
Figure 9. Distribution of token-related call number for both adversarial and benign contract samples.
4.1.6. Contract Type
Type Sample YES NO
Token benign 9245 (71.1%) 3755 (28.9%)
adversarial 0 (0%) 269 (100%)
Proxy benign 735 (5.7%) 12265 (94.3%)
adversarial 0 (0%) 269 (100%)
Selfdestruct benign 78 (0.6%) 12922 (99.4%)
adversarial 36 (13.4%) 233 (86.6%)
Table 5. Distribution of benign and adversarial contracts across different contract types.
As previously mentioned in §2.2, tokens are fundamental to DeFi. Consequently, token contracts are extensively prevalent on the blockchain, representing one of the primary types of contracts. Similarly, proxy contracts are also widely used on the blockchain, serving a different but equally vital role. They act as intermediaries to forward function calls to other contracts, allowing for flexible updates to the contract’s logic without changing the proxy contract’s address, which is important for upgradability and flexibility in smart contracts. While adversarial contracts typically do not fall into either of these types. Furthermore, the EVM provides the selfdestruct instruction, enabling a contract to delete itself from the blockchain. Some attackers have a tendency to self-destruct their contracts subsequent to the completion of the attack. As a result, we include the contract type as distinctive features.
According to our contract type identification algorithm (detailed in §4.2), in our contract samples, as shown in Table 5, none of the adversarial contracts fall into the token or proxy contract types. In contrast, over 70% of benign samples are token contracts, and proxy contracts make up more than 5%. Additionally, less than 1% of benign contracts use the selfdestruct instruction, while over 10% of adversarial contracts do so.
4.2. Feature Extraction
4.2.1. Transaction Data Retrieval
In the first step of feature extraction, we utilize an RPC (Remote Procedure Call) provider to send direct requests to nodes within the blockchain. This allows us to obtain the raw data of transactions corresponding to the contract deployment, thereby acquiring essential features such as nonce, value, input data length, and gas consumption. Additionally, the transaction sender, input data and created contract address can also be obtained, which will be utilized in subsequent phases of feature extraction.
4.2.2. Explorer Data Retrieval
Subsequently, we leverage the APIs provided by blockchain explorers, namely Etherscan and Bscscan, to easily retrieve information about the contract’s verification status.
[…]
Algorithm 1 Fund Source Labeling Algorithm
[…]
13:     end if
[…]
15:     if $n\geq 10000$ then
16:         return ’Unknown’
17:     end if
[…]
20:         return ’Unknown’
21:     end if
[…]
23:end while
[…]
Specifically, We categorize fund source addresses into four types:
  1. 1)
    Safe: Denotes fund sources with relatively secure origins, such as KYC-required centralized exchanges, which are less likely to be used by attackers.
  2. 2)
    Anonymous: Signifies fund sources with anonymous origins, including mixers and some KYC-free exchanges, which are more susceptible to be used for malicious purposes.
  3. 3)
    Bridge: Indicates funds originating from cross-chain transfers. We cannot definitively determine if their source is anonymous or not, so they are considered potentially usable for attacks.
  4. 4)
    Unknown: Addresses lacking specific labels but possessing a substantial historical transaction volume. We cannot categorize them accurately.
Figure 10. Distribution of fund source labels for both adversarial and benign contract samples.
Based on the fund source labeling algorithm, we conducted a statistical analysis for our contract samples. As depicted in Figure 10, in the benign samples, more than two-thirds of the fund source labels are identified as Safe, approximately one-quarter are labeled as Unknown, and a small portion fell into the categories of Anonymous and Bridge. Conversely, in the adversarial samples, the majority of labels are categorized as Anonymous, with a small proportion belong to the Safe category.
4.2.3. Decompilation
Finally, we employ Dedaub’s tool (Dedaub, 23) to decompile the contract bytecode, allowing us to analyze and extract the remaining features. Regardless of whether the contract is verified, to facilitate subsequent unified processing, we will perform the decompilation.
Listing 1: Abstract example of a function in contract
FUNCTION name/selector/dest public/private
    CALL name/selector
    JUMP  dest
    DELEGATECALL
    SELFDESTRUCT
Feature Model Accuracy Precision Recall F1-Score FPR
Metadata features Logistic Regression 0.9883 0.6447 0.9245 0.7597 0.0104
Decision Trees 0.9898 0.7500 0.7358 0.7429 0.0050
Random Forest 0.9906 0.7258 0.8491 0.7826 0.0065
MLP 0.9879 0.6400 0.9057 0.7500 0.0104
LightGBM 0.9910 0.7377 0.8491 0.7895 0.0062
Functionality features Logistic Regression 0.9499 0.2778 0.9434 0.4292 0.0500
Decision Trees 0.9883 0.7037 0.7170 0.7103 0.0062
Random Forest 0.9921 0.7963 0.8113 0.8037 0.0042
MLP 0.9906 0.7414 0.8113 0.7748 0.0058
LightGBM 0.9943 0.8654 0.8491 0.8571 0.0027
All features Logistic Regression 0.9894 0.6667 0.9434 0.7812 0.0096
Decision Trees 0.9959 0.9038 0.8868 0.8952 0.0019
Random Forest 0.9970 0.9091 0.9434 0.9259 0.0019
MLP 0.9936 0.7903 0.9245 0.8522 0.0050
LightGBM 0.9981 0.9286 0.9811 0.9541 0.0015
Table 6. The performance metrics for different features and models, evaluated on the complete contract samples.
Specifically, we first decompile the runtime code (deployed on the chain) of the contract to obtain the comprehensive pseudocode. Then, we employ regular expressions to extract information about implemented functions, and delve deeper into parsing the function calls within each identified function. Ultimately, the abstract representation of each function is meticulously documented, exemplified by Listing 1. Regarding public functions, Dedaub’s decompilation process strives to automatically resolve function signatures. As a result, it may either reconstruct the function names or retain the original function selectors, which are unique identifiers derived from the hash of function signatures, used by the EVM to identify and invoke functions. Conversely, for private functions, due to their invocation using jump instructions, they are represented by the destination location where the function resides.
Following this processing, we can acquire various fundamental features in both Implemented Function and Function Call categories, such as PublicFuncNumber, PrivateFuncNumber, ExternalCallNumber, InternalCallNumber, and DelegateCallNumber. Additionally, we can also determine whether the contract can self-destruct.
To glean insights into token-related calls within each function and the implementation of flashloan callbacks, we need to attempt matching each function within the contract with the functions listed in Table 4. It’s worth noting that the function resolved results by decompilation may not always be exhaustive. Therefore, we use both function names and selectors to detect token-related calls and flashloan callbacks. To further extract the MaxTokenCall and AvgTokenCall features, we analyze the execution flow of each public function, which means iteratively traversing the calls within the function, and then counting token-related calls corresponding to the involved private functions and flashloan callbacks.
Furthermore, to ascertain whether the contract is a token contract, we inspect the implemented functions to check if it fully implements all functions of a specific token ERC standard. Specifically, we implement detection for ERC-20 (ERC20, 28), ERC-721 (ERC721, 30), and ERC-1155 (ERC1155, 26) token contract standards. It’s noteworthy that other token ERC standards, such as ERC-777 (ERC777, 31) and ERC-4626 (ERC4626, 29), are backward-compatible with ERC-20. For identifying proxy contracts, despite various proxy patterns, most proxy contracts, following ERC-1967 (ERC1967, 27), typically saves the logic contract’s address in a specific storage slot. Therefore, if a contract involves a delegatecall and the contract bytecode contains the slot value, we consider it a proxy contract, by which the majority of proxy cases are covered.
4.3. Classifier Building
Once all the required features have been collected, we proceed to train a supervised learning classifier using the previously gathered datasets of benign and adversarial contracts.
Due to the limited quantity of adversarial samples, to mitigate the issue of class imbalance that can bias model predictions, we employed an oversampling technique known as ADASYN (Adaptive Synthetic Sampling). This approach synthesizes additional adversarial samples to bring the number of attack samples in the training set closer to that of benign samples.
For the fund source feature, which consists of categorical data with low cardinality, we applied one-hot encoding. Additionally, to facilitate better comparability among features of varying dimensions in numerical terms and to improve the accuracy of the classifier, we standardized the data using Z-score normalization. This process centers each feature’s values around 0 with a standard deviation of 1.
Finally, when selecting an appropriate supervised machine learning algorithm to construct the classifier, we explore various models. These models include classical approaches such as Logistic Regression and Decision Trees. We also try ensemble learning approaches, such as Random Forest based on bagging and the LightGBM based on boosting. In addition, we considered the use of a classic neural network architecture, the Multilayer Perceptron (MLP), as one of the candidate models.
5. Evaluation
In this section, we provide a comprehensive exposition of the evaluation results for the built classifier based on our dataset. We first introduce the metrics and methodologies employed for the evaluation in §5.1, then demonstrate the performance of our classifier in §5.2, elucidates the importance of features in §5.3, and finally in §5.4, we conduct a detailed analysis of the failure instances where the classifier made incorrect identifications during testing.
5.1. Metrics and Methodology
In addition to commonly used binary classification performance metrics such as accuracy, precision, recall, F1-Score, we also calculate the False Positive Rate (FPR) metric, which holds particular significance in our scenario. A high FPR implies that the classifier is more likely to incorrectly identify benign contracts as adversarial, potentially leading to significant manual audit costs.
A distinguishing aspect of our research compared to traditional supervised machine learning problems is the explicit chronological relationship between contract deployment. Different contracts deployed over time may exhibit varying characteristics and vulnerabilities, meaning that adversarial contracts may evolve as well. Additionally, our objective is to train a classifier based on existing DeFi incident data to detect potential future attacks. Consequently, traditional cross-validation techniques, which assume each data is independent, may not be suitable for our evaluation.
In response, we adopt a chronological approach to select adversarial contracts for training and testing. We use the earliest 80% of the contracts for training and reserve the latest 20% for testing. Specifically, the 53 most recent adversarial contracts, deployed between August 18, 2023, and December 21, 2023, are allocated for testing, while the remaining 216 earlier contracts are used for training.
Table 7. Comparison of performance with baseline, evaluated on the benign samples excluding token and proxy contracts.
Model Precision Recall F1-Score FPR
LR 0.7246 0.9434 0.8197 0.0315
DT 0.8491 0.8491 0.8491 0.0132
RF 0.9259 0.9434 0.9346 0.0066
MLP 0.8475 0.9434 0.8929 0.0149
LightGBM 0.9273 0.9623 0.9444 0.0066
Forta’s 0.8778 0.5520 0.6208 -
However, since the benign samples were deployed in a different time frame compared to adversarial samples, and the benign contracts are inferred based on the number of interactions, which does not unequivocally confirm a contract’s benign nature, we have adopted a different strategy for them. To maximize the accuracy of the training labels for benign contract, we choose to select the top 80% of benign contracts with the highest interaction for training. This strategy is under the assumption that contracts with more interactions are more likely to be genuinely benign. As a result, 10,400 of the most interacted benign contracts, with the lowest interaction count being approximately 1,300, are used for training. The remaining 2,600 benign contracts with fewer interactions are earmarked for testing.
5.2. Performance
Table 6 presents our test results. Overall, the LightGBM-based classifier outperforms others, demonstrating the best performance across all metrics when using all features. Its F1-Score reaches 0.9541, and it exhibits a remarkably high detection rate for adversarial samples. Out of the 53 adversarial contracts, only one remains undetected, while maintaining an exceptionally low FPR of 0.15%. Other models also perform well. Except the one based on Logistic Regression, the F1-scores of all other classifiers exceed 0.85, indicating the applicability of our approach to the dataset.
Figure 11. Top 15 LightGBM Feature Importance
To the best of our knowledge, this is the first academic work on focused the detection of adversarial contracts utilized in DeFi attack. The only similar research made public so far is by the Forta Network (FortaNetwork, 35), which also uses machine learning. However, their approach differs as it focuses solely on low-level bytecode. We used their method as a baseline for comparison. It is pertinent to mention that Forta excludes token and proxy contracts. Moreover, based on our analysis in  §4.1, it was found that adversarial contracts belonging to the token and proxy types are exceedingly rare. Therefore, to ensure a more reasonable comparison, we evaluate our classifier again by excluding token and proxy contracts from our benign contract samples. The new samples contains a total of 3,020 non-token and non-proxy benign contracts, with the most interacted 2,416 of them used for training and the remainder for testing. The results of the evaluation and comparison are presented in Table 7. It is evident that, while there is a slight decrease in classifier performance compared to using all benign samples, our classifiers still significantly outperform Forta’s most recently published model (released on February 6, 2023) across all performance metrics. This highlights the superiority of our method.
Table 8. Failure instances of the lightGBM-based classifier with a subset of important features. The first contract represents a false negative located on BSC, while the remaining four contracts correspond to false positives on the Ethereum.
Contract Address
Nonce
Fund Source
InputDataLength
GasUsed
Verified
PublicFuncNumber
FlashloanNumber
TokenCallNumber
AvgTokenCall
SelfDestructive
0x0f41f9
3
Anonymous
2616
337467
False
1
0
0
0
False
0xdac3d4
102
Bridge
21060
2195797
False
6
1
10
1.2
True
0x1d2cbc
52
Anonymous
26762
2956727
False
16
0
11
0.75
True
0x70d2f0
0
Anonymous
19336
2099607
False
13
0
12
1.08
True
0xfd594e
1
Anonymous
3306
421261
False
3
0
2
0.67
False
5.3. Feature Importance
To assess the impact of different feature categories, we divided the features into two major classes: one focused on the contract metadata, including Creator Attribute, Transaction Data, and Verification Status, the other related to the contract functionality, including Implemented Function, Function Call, and Contract Type. We conducted tests using only one feature class at a time, and the results are presented in Table 6.
Clearly, when compared to using all features, training classifiers with a single class of features leads to a certain degree of performance degradation for all models. This indicates that each category of features plays a valuable role in the detection of adversarial contracts. Moreover, it’s worth noting that the performance varies across different models when utilizing the two feature classes separately. Random Forest, MLP, and LightGBM models tend to perform better when using functionality features, while the other two models exhibit better performance with metadata features. This observation suggests that the importance of different features varies depending on the model used.
Figure 11 illustrates the top 15 important features of the lightGBM-based classifier when all features are used. It is evident that various feature categories in this classifier contribute significantly to its performance.
5.4. Failure Analysis
In this section, we analyze the failure instances of the LightGBM-based classifier, where the classifier made incorrect identifications during testing. Specifically, there is one false negative and four false positives. Table 8 presents these failure contracts along with a subset of their important features.
5.4.1. False Negative
The adversarial contract with address 0x0f41f9 stands out as the sole false negative instance, erroneously classified as benign by the classifier. While it exhibits some characteristics typically associated with adversarial contracts, such as an Anonymous fund source, lack of verification, a single public function, and relatively low input data and gas consumption, it notably lacks the implementation of the flashloan callback and does not involve token-related calls. These deviations from the typical adversarial contract might have contributed to the misclassification by the classifier.
This incident occurred on September 29, 2023, resulting in approximately 4,000 USD in losses. Upon analyzing the victim contract and the adversarial transaction, it becomes apparent that the vulnerability exploited in this attack was an access control flaw that led to arbitrary external call. In the adversarial contract, the attack was completed with only three external calls. (1) It first invoked the update function within the victim contract. This function could modify critical variables but lack proper access control, enabling anyone to call it. (2) Next, the adversarial contract exploited the functionCallWithValue function of the victim contract, transferring its native tokens to adversarial contract specified in the previously called update function. (3) Finally, the acquired tokens were transferred to the attacker. This attack was relatively simple, with losses directly in native tokens and no involvement of token-related calls, suggesting that our detection method may have limitations when confronted with adversarial contracts that employ straightforward attack flow.
5.4.2. False Positive
The four false positive instances share certain characteristics resembling typical adversarial contracts, such as non-Safe fund source labels, unverified status, the usage of token-related calls, and the ability to self-destruct. These particular features might have led the classifier to identify them as adversarial contracts.
Arbitrage. The contract 0xdac3d4 was found to be involved in arbitrage activities through an analysis of its transaction history. Arbitrage involves profiting from price differences in tokens across different platforms. Specifically, this contract was used for the NFT arbitrage. Taking its first transaction after deployment, transaction 0x25637d, as an example, the contract initially purchased an NFT on the Blur (blur, 9) for 0.3 Ether and immediately sold it on the Sudoswap (sudoswap, 56) for approximately 0.314 Ether, thereby profiting from the price difference. The entire process was executed within a single transaction using the contract and leveraged the private service of flashbots to mitigate the risk of failure.
SniperBots. The contract 0x1d2cbc has already self-destructed. Upon analyzing its transaction history, it was observed that nearly all of its transactions involved purchasing ERC-20 tokens with Ether. Furthermore, the purchase transaction and the transaction for creating the liquidity pool corresponding to the target token both occurred within the same block. Consequently, this contract was actually used by sniper bots (cernera2023token, 14). These bots engage in buying tokens on liquidity pools as soon as they are listed, allowing them to ”snipe” new tokens at their initial price before anyone else, with the intention of selling them at a profit once their value increases. The contract 0x70d2f0 was also utilized by similar sniper bots.
Phishing. The contract 0xfd594e has relatively simple code. An analysis of its transaction history reveals a consistent pattern in all transactions, which is phishing activity in fact. Specifically, it was employed for zero-value transfer phishing. In this phishing method, scammers begin by generating fake addresses with characters matching the first and last few characters of the victim’s recent interaction address. Subsequently, they employ the transferFrom function to perform zero-value transfer to the victim’s address. This action results in the appearance of the fake address in the victim’s transaction history, misleading users with the intent of causing them to inadvertently send tokens to this fraudulent address, thereby stealing the victim’s tokens.
These false positives suggest that contracts used for arbitrage, as well as those employed in sniper bots, phishing, and other suspicious activities aimed at financial gain, bear certain similarities to adversarial contracts involved in DeFi attacks. These may represent the primary contract types responsible for the false positives in our detection methodology.
6. Limitations and Future Work
In this section, we discuss the limitation as well as the possible future work.
Detection Scope. Our current method has certain limitations in the scope of detection. In terms of feature selection, we rely on existing attack patterns, with a specific focus on DeFi attacks involving financial losses, tailored to prevalent attack scenarios. Consequently, we may miss attacks that do not result in financial losses and certain adversarial contracts with novel attack vectors. Furthermore, in the cat-and-mouse game of DeFi attack and defense, we have observed that some DeFi attackers, similar to front-running bots, encapsulate function calls related to attacks within the input data of adversarial transactions. As the attacking logic employing this approach is not included in the contract, it also poses difficulties for detection.
Performance. As analyzed in section §5.4, our classifier struggles to correctly categorize adversarial contracts with overly simplistic attacking logic, as well as contracts designed for suspicious activities aimed at financial gain, which significantly impacts the accuracy of our detection system. Additionally, the current feature extraction process relies on RPC providers and blockchain explorer APIs. While these components can be replaced by running our custom nodes, as of writing, there is a lack of robust open-source decompilation tools. Solely relying on Dedaub’s online decompilation service results in slow response times, making real-time detection impractical.
In the future, we will continue to refine our contract dataset, which involves not only supplementing the adversarial contracts but also gathering a more diverse range of benign contracts from a broader time span. Additionally, we plan to delve deeper into the nuanced differences between contracts utilized in suspicious activities such as arbitrage, sniper bots, phishing, and contracts implicated in DeFi attacks, helping us extract more representative features to reduce false positives while maintaining efficient detection capability for adversarial contracts. We are also contemplating a more granular classification of adversarial contracts based on specific attacking logic they employ, enabling us to identify the vulnerabilities that the attackers intent to exploit in the victim contract. Furthermore, we recognize the potential benefits of developing our own decompilation module. This would not only improve the real-time performance of feature extraction but also allow for more comprehensive features from bytecode.
7. Related Work
In this section, we introduce two categories of closely-related work.
7.1. Adversarial Transaction Detection
Some researches (chen2020soda, 17, 38, 53, 69) modify EVM to obtain more detailed transaction execution information and provide scalable interfaces for online detection of various types of adversarial transactions. Sereum (rodler2018sereum, 53) uses taint analysis to monitor runtime data flows during the execution of transactions to detect reentrancy attacks. SODA (chen2020soda, 17) is a generic online detection framework that can be used to detect various adversarial transactions based on predefined rules. EthScope (wu2022time, 69) is a scalable Ethereum attack detection framework which stores historical data into an cluster database and incorporates dynamic taint analysis into the client to analyze transactions.
On the other hand, some researches (gai2023blockchain, 37, 49, 64, 70, 72, 76) achieve adversarial transaction detection through forensic analysis. TxSpector (zhang2020txspector, 72) is a logic-driven framework designed to analyze Ethereum transactions, identifying attacks according to user-defined detection rules by encoding the transaction trace into logical relationships. DeFiScanner (wang2022defiscanner, 64) is the first deep-learning-based attack detection system, capable of identifying adversarial transactions based on features from both the account and transaction levels. BLOCKGPT (gai2023blockchain, 37) applies large language model into abnormal transaction detection and demonstrates strong performance in detecting adversarial transactions. Qin et al. (qin2023towards, 49) propose an Execution Property Graph and a graph traversal technique which can swiftly identify potential adversarial transactions efficiently.
7.2. Abnormal Contract Detection
Forta‘s (FortaNetwork, 35) method combines Natural Language Processing (NLP) techniques to identify adversarial contracts at the bytecode level, and apart from it, there is no existing work that specifically focuses on the detection of adversarial contracts. The majority of abnormal contract detection focuses on Ponzi schemes and honeypots. They share similarities in methods and can be categorized into code similarity detection (bartoletti2020dissecting, 7, 57), machine learning (camino2020data, 12, 18, 19, 39, 40, 41, 74), and symbolic execution techniques (chen2021sadponzi, 20, 61).
Ponzi scheme. The traditional Ponzi scheme is a financial fraud that promises high returns to early investors by using funds from new investors, ultimately collapsing when there are not enough new investors, leaving most participants with losses. The emergence of smart contracts has enabled the replication of Ponzi schemes on blockchain platforms. Chen et al. (chen2018detecting, 19) focused on extracting features based on the contract bytecode and the transaction information of the contract. Ji et al. (ji2023ponzi, 41) extracted features based on the control flow graph of the contract. Zheng et al. (zheng2023securing, 74) constructed a larger dataset of Ponzi scheme than previous work (chen2018detecting, 19) and proposed a Multi-view Cascade Ensemble method without relying on contract interaction information.
Honeypots. Honeypots are a distinct type of contracts designed to entice victims into traps, employing seemingly vulnerable structures to conceal risks rather than relying on traditional deceptive methods. Torres et al. (torres2019art, 61) developed a taxonomy of honeypot techniques and built a tool called HoneyBadger that employs symbolic execution and well defined heuristics to expose honeypots. Subsequently, Camino et al. (camino2020data, 12) introduced a data science approach based on the transaction behavior of contracts. Chen et al. (chen2020honeypot, 18) constructed a classification model based on the opcode of contracts.
8. Conclusion
In this paper, we propose a new direction for detecting DeFi attacks, i.e., detecting adversarial contracts instead of adversarial transactions. We build the first dataset for adversarial contracts. Based on this dataset, we found that the proportion of adversarial transactions sent using privacy services has exhibited a significant upward trend, thereby evading traditional adversarial transaction detection methods. Through the analysis of DeFi attack patterns, we extract six categories of features across multiple dimensions, such as attributes of the contract creator, functions implemented in the contract, and types of contracts. Utilizing these features, we build a machine learning classifier that can effectively distinguish adversarial contracts from benign ones. The evaluation show that our method performs exceptionally well. The F1-Score for LightGBM-based classifier is 0.9541, with a remarkably low false positive rate of only 0.15%. By analyzing cases where the classifier made incorrect identifications, we point out contracts that could be confused with adversarial contracts used in DeFi attacks.